OpenStudy (lichking):

Medal to the best answer. At a party, everyone shook hands with everybody else. There were 66 handshakes. How many people were at the party?

OpenStudy (anonymous):

I've heard of this problem before, I think you'll need quadratics. :/ Sorry, can' help here as I haven't learnt it yet. Good luck!

OpenStudy (mathmate):

You're probably doing graph theory. If you are, think of degree of complete graphs.

OpenStudy (anonymous):

\[\left(\begin{matrix}n \\ 2\end{matrix}\right)-2=66\] then \[\left(\begin{matrix}n \\ 2\end{matrix}\right)=64\] implies that n=12

OpenStudy (lichking):

12 is correct

OpenStudy (lichking):

but i need an explpicit method

OpenStudy (anonymous):

I regret sir, i only know this method

OpenStudy (mathmate):

Let there be n people at the party. Each person shook hands with (n-1) others. Total of n(n-1) handshakes, but we double counted because it takes two people to make a handshake. So we end up with n(n-1)/2=66 Solve for n.